The perimeter of a rectangle is 70 inches. If the length of the rectangle is eight inches more than two times the width, find the dimensions.
Answer by TimothyLamb(4379)

x = length (inches)
y = width (inches)
---
2x + 2y = 70
x = 2y + 8
---
put the system of linear equations into standard form
---
2x + 2y = 70
x - 2y = 8
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 26
y = 9
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---